OpenStudy (anonymous):

Verify the identity. Show your work. (1 + tan2u)(1 - sin2u) = 1 my answer > cos^2x+sin^2x=1

OpenStudy (anonymous):

but my teacher said 6. You can re-write each factor as another trig function squared. 1+tan^2u=(standard identity)? and 1-sin^2u=(standard identity)? Once you do that, the next steps are clear. Please refer to section 5.1 of your text for examples and sample problems.

OpenStudy (anonymous):

@phi @Vocaloid

OpenStudy (freckles):

what does 6 mean?

OpenStudy (freckles):

also do you mean tan^2(u) and sin^2(u) because I see tan(2u) and sin(2u)

OpenStudy (freckles):

Also are you asking us to refer to section 5.1 of your book?

OpenStudy (freckles):

I don't see how that is possible.

OpenStudy (loser66):

I am with @freckles It must be \((1+tan^2u)(1-sin^2u) =1\)

OpenStudy (anonymous):

idk yes my book 6 the number of the problem im on and idk what she means '

Nnesha (nnesha):

well you are right cos^2+ sin^2 =1 right ! BUT you need to prove( 1+tan^2 ) (1-sin^2i) equal to 1

OpenStudy (anonymous):

sohow do i do that

Nnesha (nnesha):

\[\huge\rm (1+\tan^2u)(1-\sin^2u)=1\] show that R.H.S=L.H.S are u familiar with the trig identities you have to apply that :=)

OpenStudy (freckles):

\[\cos^2(x)+\sin^2(x)=1 \\ \text{ subtract } \sin^2(x) \text{ on both sides } \\ \cos^2(x)=1-\sin^2(x) \\ \text{ hint: now divide both sides of } \cos^2(x)+\sin^2(x)=1 \\ \text{ by } \cos^2(x)\]

OpenStudy (freckles):

other than that I can't refer to any thing in your book because I don't think I have your book

OpenStudy (freckles):

but i can help you prove the identity above

OpenStudy (anonymous):

lol i just need help with it i seriouly dont know what else to do

OpenStudy (freckles):

try to see if you can follow what I said above

OpenStudy (freckles):

divide both sides by cos^2(x) of the identity I spoke about just now and tell me what you have

OpenStudy (anonymous):

i have sin= 1 right idk im horrible at math

OpenStudy (freckles):

Did you try to divide both sides of cos^2(x)+sin^2(x)=1 by cos^2(x)?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

so do you know what cos^2(x)/cos^2(x)=? or what sin^2(x)/cos^2(x)=? or what 1/cos^2(x)=?

OpenStudy (anonymous):

how did you gwt all that

OpenStudy (freckles):

Well I asked you to divide both sides of cos^2(x)+sin^2(x)=1 by cos^2(x)

OpenStudy (freckles):

\[\frac{\cos^2(x)+\sin^2(x)}{\cos^2(x)}=\frac{1}{cos^2(x)}\]

OpenStudy (anonymous):

ohhh i did it way wrong

OpenStudy (freckles):

\[\frac{\cos^2(x)}{\cos^2(x)}+\frac{\sin^2(x)}{\cos^2(x)}=\frac{1}{\cos^2(x)}\]

OpenStudy (freckles):

cos^2(x)/cos^2(x)=1 sin^2(x)/cos^2(x)=tan^2(x) 1/cos^2(x)=sec^2(x)`

OpenStudy (freckles):

though you could just leave it as 1+tan^2(x)=1/cos^2(x) if you want

OpenStudy (freckles):

now plug in both of those results we got from the Pythagorean identity

OpenStudy (loser66):

\(1+tan^2u = sec^2 u =\dfrac{1}{cos^2u}\\1-sin^2 u = cos^2u\\hence,(1+tan^2u)(1-sin^2u)= \dfrac{1}{\cancel{cos^2u}}*\cancel{cos^2u}=1\)